OpenStudy (anonymous):

Questions you didn't probably know could be answered with math.(Already created formulas)

OpenStudy (anonymous):

WHY WE PROCRASTINATE (Delay tasks)

OpenStudy (anonymous):

The Formula: \[\Huge U=\frac{ EV }{ ID }\]

OpenStudy (anonymous):

Here, U is the desirability of the task, E is the confidence that you will succeed at it, V is the value of completing it, I is the task's immediacy or availability and D is your sensitivity to delay. Consider for example the prospect of studying for a big exam. We'll put your confidence at 5 because you haven't actually been to any lectures this semester. The value of the task can be 5, too, because reading about third-year accounting rips your soul through your nasal passages. Immediacy can be set to 2 because you know you're going to have to invest the entire night in it, and delay sensitivity we'll put at 10 because even the drapes have a good chance of distracting you at this point. Crunch the numbers through Steel's formula and we see the desirability of this task is about 1.25.

OpenStudy (anonymous):

2-What Makes a Movie Scary:-

OpenStudy (anonymous):

The Formula: \[ (es+u+cs+t)^{2} + s + \frac{ tl+f }{ 2 }+ \frac{ a+dr+fs }{ n }+sinx - 1 \]

OpenStudy (anonymous):

In this ridiculously complicated formula: es is escalating music; u is the unknown; cs is chase scenes; t is the sense of being trapped; a is the character being alone; dr is how dark the film is; fs is the film setting; tl stands for true life; f stands for fantasy; n is for number of people; sin is blood and guts and s is shock.

OpenStudy (anonymous):

3 . The Best Time for a Marriage Proposal

OpenStudy (anonymous):

Don't worry, you don't actually have to calculate that intimidating algebraic mess. Professor Tony Dooley from the University of New South Wales School of Mathematics and Statistics, who came up with the formula, assures us that the answer is 0.368, and we all have to take his word for it. The only thing you need to do is pick the youngest age you would want to get married, subtract that from the oldest age and multiply that by 0.368. Then you add that number to your youngest age, and voila. For example, if you start thinking of settling down at 21, but you don't mind being single until you're 30, then: (30-21) x 0.368 = 3.312 + 21 = 24 and three months, which is when you should consider popping the question.

OpenStudy (anonymous):

4. How to Ride a Bike

OpenStudy (anonymous):

1.The Secret to Happiness

OpenStudy (anonymous):

\[\huge Happiness=P+(5*E)+(3*H)\]

OpenStudy (anonymous):

Here, P stands for personal characteristics, E for existence (health, financial stability and friendships) and H is higher order, and covers self-esteem, confidence, ambitions and sense of humor

OpenStudy (anonymous):

@ganeshie8 @AkashdeepDeb @nikato @Kainui @Miracrown @iambatman @hartnn @AravindG @nincompoop @Mashy @BSwan @chmvijay @dan815 @Directrix @ikram002p @Hero @jim_thompson5910 @LolWolf @math_genius123 @mathmate @mathslover @matricked @OrangeMaster @PeterPan @perl @sidsiddhartha @sidsiddhartha @vishweshshrimali5 @Vincent-Lyon.Fr @Zale101 @tkhunny @thomaster @satellite73 @SolomonZelman @ranga

OpenStudy (anonymous):

@Preetha

OpenStudy (akashdeepdeb):

Try sending this to the Noble foundation.

OpenStudy (anonymous):

@shrutipande9

OpenStudy (anonymous):

there are some abstract values , which you might think can't be measured , like personal charachteristics , it can be measured using a standard scale known worlwide and thus can be input into the formulae

OpenStudy (here_to_help15):

I am not understanding what you need help with?

OpenStudy (ikram002p):

loved it :) + i know about it before but th for sharing

OpenStudy (camerondoherty):

Nice observations! Theyre really cool